The pressure of a gas changes linearly with volume from 10 kPa, 200 cc to 50 kPa, 50 cc. (a) Calculate the work done by the gas. (b) If no heat is supplied or extracted from the gas, what is the change in the internal energy of the gas?

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. P base 1 = 10 kpa = 10 × 10^3 pa. P base 2 = 50 × 10^3 pa. v base 1 = 200 cc. v base 2 = 50 cc (i) Work done on the gas = ½ (10+50) x 10^3 x (50-200) x 10 ^-6 = - 4.5 J (ii) dQ = 0 ⇒ 0 = du + dw ⇒ du = – dw = 4.5 J